### v1

# LLM-JEPA: Large Language Models Meet Joint Embedding Predictive Architectures

Abstract
Large Language Model (LLM) pretraining, finetuning, and evaluation rely on input-space reconstruction and generative capabilities. Yet, it has been observed in vision that embedding-space training objectives, e.g., with Joint Embedding Predictive Architectures (JEPAs), are far superior to their input-space counterpart. That mismatch in how training is achieved between language and vision opens up a natural question: can language training methods learn a few tricks from the vision ones? The lack of JEPA-style LLM is a testimony of the challenge in designing such objectives for language. In this work, we propose a first step in that direction where we develop LLM-JEPA, a JEPA based solution for LLMs applicable both to finetuning and pretraining. Thus far, LLM-JEPA is able to outperform the standard LLM training objectives by a significant margin across models, all while being robust to overfiting. Those findings are observed across numerous datasets (NL-RX, GSM8K, Spider, RottenTomatoes) and various models from the Llama3, OpenELM, Gemma2 and Olmo families. Code: https://github.com/rbalestr-lab/llm-jepa.
Figure 1: LLM-JEPA produces strong fine-tuned models across datasets and models.
1 Introduction
The research landscape around representation learning has been increasingly divided into two camps: (i) generative or reconstruction-based methods brown2020language ; chowdhery2023palm ; he2022masked ; lecun2022path , and (ii) reconstruction-free Joint Embedding Predictive Architectures (JEPAs) assran2023self ; baevski2022data2vec ; bardes2024revisiting . While the former is self-explanatory, the latter learns a representation by ensuring that different views, e.g., pictures of a same building at different time of day, can be predicted from each other, all while preventing a collapse of the embeddings. By moving away from input-space objectives, JEPAs training benefits from less biases littwin2024jepa , at the cost of potential dimensional collapse of their representation jing2021understanding ; kenneweg2025jepa . That divide has been well studied in vision, where it was found that JEPAs offer multiple provable benefits when it comes to knowledge discovery for perception tasks. In the realm of Natural Language Processing however, reconstruction-based methods remain predominant. In fact, today’s Large Language Models are mostly judged from their ability to generate samples and answers in input space in text form–making it challenging to leverage JEPA objectives.
Yet, LLMs’ task also involve perception and reasoning where JEPA is known to be preferable. It thus seems crucial to adapt JEPA solutions to LLMs in the hope to showcase the same benefits as witnessed in vision. This first step is exactly what we present in this study. We propose to improve the representation quality of LLMs by leveraging a novel objective combining both the original reconstruction based loss–with an additional JEPA objective. To do so, we focus first on tasks and datasets that are inherently suited for JEPA objectives: the ones providing multiple views of the same underlying knowledge. One typical example is a git issue and the corresponding code diff (fig.˜2) jimenez2024swebench . The two samples are two views–one being plain English and one being in code–of the same underlying functionality. Let’s use that particular example to highlight our core contribution:
Viewing the (text,code) pairs as views of the same underlying knowledge enables JEPA objectives to be utilized with LLMs, complementing the standard text $\rightarrow$ code generative task.
We strongly emphasize that being able to obtain non-trivial views, such as described above, is crucial to the success of JEPA objectives. While we restrict ourselves to datasets offering those non-trivial views, developing a mechanism akin to data-augmentation in vision would enable JEPA objectives to be used on any dataset. Nonetheless, we believe that our proposed solution–coined LLM-JEPA–and empirical study will serve as a first step towards more JEPA-centric LLM pretraining and finetuning. We summarize our contributions below:
  • Novel JEPA-based training objective: We present the first JEPA-based training objective for LLMs operating in embedding space and with different views–perfectly following vision-based JEPAs without sacrificing the generative capabilities of LLMs
  • Improved SOTA: We empirically validate our formulation in various finetuning settings, where we obtain improvements over standard LLM finetuning solutions. We also explore pretraining scenarios showing encouraging results of LLM-JEPA
  • Extensive empirical validation: on various model family (llama, gemma, apple/openelm, allenai/olmo), dataset (NL-RX, GSM8K, Spider, RottenTomatoes), and size.
Natural Language to Regular Expression
Natural Language to SQL
Figure 2: Left: JEPA applied to NLP tasks that has TextText and $Code$, where TextText and CodeCode are naturally two views of the same thing. Right: (top): An illustration of the NL-RX-SYNTH dataset, where each sample consists of a description of the regular expression in natural language (TextText) and the regular expression itself (CodeCode). (bottom): The Spider dataset, where $Text$ is the database ID and description of the SQL query and $Code$ is the SQL query itself.
2 JEPA-LLM: Improving LLMs’ Reasoning and Generative Capabilities
The first section˜2.1 provides minimal background around next-token prediction LLM objectives, used as part of the proposed LLM-JEPA loss (section˜2.2). Empirical validation will then be provided in section˜2.3 demonstrating clear finetuning and pretraining benefits.
2.1 Primer on Large Language Models
Contemporary LLMs are mostly built from the same core principles: stacking numerous layers of nonlinear operations and skip-connections–known as Transformers. While subtleties may differ, e.g., about positional embeddings, initialization, normalization, the main driver of performance remains the availability of high quality dataset during the pretraining stage. The training objective in itself has also been standardize throughout methods: autoregressive token-space reconstruction. Let’s first denote by $\mathcal{L}_{\rm LLM}$ the typical LLM objective used for the specific task and dataset at hand. In most cases, this will be a cross-entropy loss between the predicted tokens and the ground-truth token to reconstruction. We note that our LLM-JEPA construction is agnostic of $\mathcal{L}_{\rm LLM}$ hence making our method general to numerous scenarios.
$\mathcal{L}_{\rm LLM}({\rm Text}_{1:L-1},{\rm Text}_{L})={\rm XEnt}\left({\rm Classifier}\left({\rm Enc}({\rm Text}_{1:L-1})\right),{\rm Text}_{L}\right),$ (1)
where $\rm Classifier$} predicts the logits of the next token ${\rm Text}_{L}$ given the past tokens ${\rm Text}_{1:L-1}$. Computation of eq.˜1 is done at once over $L$ through causal autoregression. Different stages and tasks may vary the input and output of the loss.
2.2 The LLM-JEPA Objective
Throughout this section, we will use $\rm Text$} and $\rm Code$} as concrete examples of having different views of the same underlying knowledge. It should be clear to the reader that our proposed LLM-JEPA objective handles different types of views similarly.
The construction of our LLM-JEPA objective relies on two principles. First, we must preserve the generative capabilities of LLMs and we therefore start with the $\mathcal{L}_{\rm LLM}$ from eq.˜1. Second, we aim to improve the abstraction capabilities of LLMs using the joint embedding prediction task. On top of $\mathcal{L}_{\rm LLM}$, we then propose to add the well-established JEPA objective leading to the complete loss $\mathcal{L}$ defined as
$\mathcal{L}_{\rm LLM-JEPA}=\underbrace{\sum_{\ell=2}^{L}\mathcal{L}_{\rm LLM}({\rm Text}_{1:\ell-1},{\rm Text}_{\ell})}_{\text{generative capabilities (LLM)}}+\lambda\times\underbrace{d({\rm Pred}({\rm Enc}({\rm Text})),{\rm Enc}({\rm Code}))}_{\text{abstraction capabilities (JEPA)}},$ (2)
where $\lambda\geq 0$ is an hyperparameter balancing the contribution of the two terms, Pred and Enc are the predictor and encoder networks respectively, and $d$ is a metric of choice, e.g., the $\ell_{2}$ distance. Let’s now precisely describe each of those components.
The encoder. We use the hidden_state of the last token from the last layer as the embedding of an input sequence–as commonly done for LLM probing. Practically, we can not produce ${\rm Enc}({\rm Text})$ and ${\rm Enc}({\rm Code})$ through a single forward pass. For example, passing the concatenation of ${\rm Text},{\rm Code}]$ would require meddling with the self attention to avoid cross-view interaction which would be efficient but specific to each LLM architecture. Instead, we propose to get the encoding through two additional forward passes: one for $\rm Text$}, and one for $\rm Code$}. This incurs additional costs during training–but not during inference–see section˜3 for further discussions.
The metric. When it comes comparing embeddings, it is now widely accepted in vision to leverage the cosine similarity. We thus propose to do the same for LLM-JEPA.
The predictor. We leverage the auto-regressive nature of LLM and their internal self-attention to define a tied-weights predictor. By introducing a special token [PRED] at the end of a given input, we allow for further nonlinear processing of the input hereby producing $Pred(\cdot)$ at the final embedding of the last layer. By reusing the internal weights of the LLM for the prediction task, we greatly reduce the training overhead and architectural design choices. Practically, we append $k\in\{0,\dots,K\}$ predictor tokens to an input prompt and use the embedding of the last predictor token to be $Pred(Enc(\cdot))$. When $k=0$, the predictor is trivial, i.e., $Pred(x)=x$.
Relation to Previous Work. Because loss functions such as $\mathcal{L}_{\rm LLM}$ (input space reconstruction since tokens are lossless compression of the original prompts) have been shown to be sub-optimal in vision, a few LLM variations have started to employ embedding space regularizers and training objectives barrault2024large ; wang2025reversal . Current solution however rely on intricate structural constraints of the embedding space, e.g., hierarchical organization and cluster, and thus fall out of the JEPA scope. We also note that our interpretation of views when it comes to LLM datasets, e.g., (text issue, code diff), is something that has been leveraged as part of the LLM finetuning solutions–by learning to generate one from the other–without a JEPA-style loss. This includes natural language to regular expression translation (locascio2016neural, ; ye2020sketch, ; zhong2018semregex, ), natural language to SQL parsing (guo2019towards, ; iyer2017learning, ; li2023resdsql, ; wang2019rat, ; yu2018spider, ) and the more recent issue descriptions to code diffs (cabrera2021commit2vec, ; hoang2020cc2vec, ; tian2020evaluating, ; zhou2023ccbert, ). More intricate examples involve text-based problem solving and their counterpart program induction (amini2019mathqa, ; cobbe2021training, ; hendrycks2021measuring, ; ling2017program, ).
2.3 Empirical Validation: LLM-JEPAs Outperforms LLMs
The JEPA loss is not implicitly minimized by $\mathcal{L}_{\rm LLM}$. The very first observation we want to make, provided in fig.˜4 lies in observing that minimizing $\mathcal{L}_{\rm LLM}$ does not implicitly minimize $\mathcal{L}_{\rm JEPA}$–indicating that it is required to add that term during training.
LLM-JEPA Improves Finetuning. We run experiments across multiple pretrained LLMs (Llama-3.2-1B-Instruct (llama3, ), gemma-2-2b-it (team2024gemma, ), OpenELM-1_1B-Instruct (mehta2024openelm, ), and OLMo-2-0425-1B-Instruct (olmo20242, )) with various datasets (NL-RX-SYNTH, NL-RX-TURK (locascio2016neural, ), GSM8K (cobbe2021training, ), Spider (yu2018spider, )). For a given (model,dataset) case, search for the best learning rate $lr\in\{1e-5,2e-5,4e-5,8e-5\}$ based on the best possible accuracy of $\mathcal{L}_{\rm LLM}$ after $4$ epochs. Then we tune the hyperparameter specific to $\mathcal{L}_{\rm LLM-JEPA}$, $k$ and $\lambda$ in a two dimensional grid defined by $(k,\lambda)\in\{0,1,2,3,4\}\times\{0.5,1,2,4\}$ (fig.˜3 and table˜5). For both NL-RX-SYNTH and NL-RX-TURK, accuracy is exact match of the generated regular expression; for GSM8K, accuracy is exact match of the final result; and for Spider, accuracy is exact match of the execution result of the generated query. We provide results demonstrating how LLM-JEPA improves performances (fig.˜1 left) across models (table˜8), datasets (table˜9), training time (fig.˜1 right and fig.˜5), and sizes (table˜11). Examples of inputs and targets along with models’ predictions and error analysis are provided in tables˜6 and 7. The improved performance of LLM-JEPA holds across LoRA ranks as shown in table˜3. We also provide evidence that LLM-JEPA induces an approximately linear transformation from $\operatorname{Enc}(\text{Text})$ to $\operatorname{Enc}(\text{Code})$ (figs.˜6, 7 and 10).
LLM-JEPA Improves Pretraining. We pretrain Llama-3.2-1B-Instruct from randomly initialized weights on NL-RX-SYNTH dataset, a prediction is valid as long as it starts with the ground truth. We obtain that LLM-JEPA also improves the quality of the learned representation, as shown in table˜1. We also conduct another pretraining experiment on cestwc/paraphrase containing groups of 5 paraphrases. We employ the paraphrases within a same group for the JEPA loss. Once the model is pretrained (4 epochs), we do finetuning evaluation on rotten_tomatoes (1 epoch). We demonstrate how JEPA pretraining improves the downstream performance post-finetuning in table˜4. Note that finetuning does not employ the JEPA loss–hence showing the benefit of JEPA at pretraining stage. Lastly, we provide in table˜2 generated samples demonstrating that JEPA pretraining does maintain the generative capabilities of the model when prompted with the first few tokens in the cestwc/paraphrase dataset.
Table 1: Pretraining accuracy on dataset NL-RX-SYNTH by Next Token Prediction ($\mathcal{L}_{\rm LLM}$) loss vs. $\mathcal{L}_{\rm LLM-JEPA}$ loss (our method). We inherit the best configuration from fine-tuning. Each case runs five times. Average accuracy and standard deviation are reported. We also report $p$-value of paired, single-tailed $t$-Test.
Model Method Accuracy (%) $\uparrow$ $p$-value $\downarrow$ Config
Llama-3.2-1B-Instruct $\mathcal{L}_{\rm LLM}$ $54.38\pm 1.70$ 2.94e42.94e-4 lr=8e5lr=8e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $60.59\pm 1.01$ $\lambda=2,k=3$, same lrlr
3 Conclusion and Future Work
We introduced an alternative training objective for LLMs leveraging JEPAs. Our formulation is an exact replicate of the JEPA objective extensively used in vision–but that hadn’t been adapted to language yet. Crucially, our proposed LLM-JEPA maintains the generative capabilities of LLMs while improving their abstract prompt representation as empirically validated across datasets and models. While our experiments mostly focus on finetuning, preliminary pretraining experiment are promising which we plan to scale and more thoroughly test in future work. Regarding the limitations of LLM-JEPA, the main current bottleneck is the 3-fold compute cost during training required to obtain the representations of the views. We plan to explore possible mitigation that would mask the self-attention matrix and allow for our LLM-JEPA loss to be evaluated within a single forward pass through the LLM.
Appendix A Appendix
Table 2: Generated samples by model pretrained by paraphrase dataset. The pretrained model is not good at terminating sentence. prompt and generation
Ground Truth vs. Generation
Ground Truth A garden of flowers and a bench stating "City of London."
Generation A garden of flowers and a vase with a flower in it………….
Ground Truth A person that is riding on a horse in a grass field.
Generation A person that is riding in a field……………..
Ground Truth A man is riding a horse in a field.
Generation A man is riding a horse in a field…………….
Ground Truth There are two birds standing on top of a building
Generation There are two birds standing on a rock……………..
Ground Truth Two hawks sit on top of a roof spire.
Generation Two hawks sit on top of a wooden bench…………….
Ground Truth .A young woman serving herself at a cookout.
Generation .A young woman serving herself in a kitchen……………..
Ground Truth 2 bowls of fruit sit on a table.
Generation 2 bowls of fruit sit on a table……………..
Ground Truth A wooden bench written ’CITY OF LONDON’ at the park
Generation A wooden bench written ’CITY and a tree……………..
A.1 Faster LoRA Convergence
Table˜3 demonstrates that LoRA fine-tuning with $\mathcal{L}_{\rm LLM-JEPA}$ loss not only achieves substantially higher accuracy than using $\mathcal{L}_{\rm LLM}$ alone, but also converges more quickly. Notably, at a LoRA rank of $512$, our method already reaches accuracy comparable to full fine-tuning, whereas LoRA with only $\mathcal{L}_{\rm LLM}$ still exhibits a clear performance gap.
Table 3: Fine-tuning accuracy on dataset NL-RX-SYNTH, LoRA vs. full fine-tuning, both by $\mathcal{L}_{\rm LLM}$ loss and $\mathcal{L}_{\rm LLM-JEPA}$ loss (our method). Configuration is $lr=2e-5,\lambda=1,k=1$. Each cell runs five times. Average accuracy and standard deviation are reported. At every LoRA rank, $\mathcal{L}_{\rm LLM-JEPA}$ (ours) has better accuracy. At LoRA rank 512 (22.59% trainable parameters), $\mathcal{L}_{\rm LLM-JEPA}$ (ours) achieves same accuracy as full fine-tuning, but $\mathcal{L}_{\rm LLM}$ still has a significant gap from full fine-tuning.
LoRA Rank Method Accuracy (%) $\uparrow$
$32$ $\mathcal{L}_{\rm LLM}$ $6.09\pm 0.55$
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $7.45\pm 1.87$
$64$ $\mathcal{L}_{\rm LLM}$ $21.09\pm 1.90$
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $32.46\pm 1.26$
$128$ $\mathcal{L}_{\rm LLM}$ $34.21\pm 2.82$
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $48.45\pm 3.66$
$256$ $\mathcal{L}_{\rm LLM}$ $45.57\pm 4.52$
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $60.80\pm 2.31$
$512$ $\mathcal{L}_{\rm LLM}$ $50.18\pm 5.15$
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $72.41\pm 2.94$
Full $\mathcal{L}_{\rm LLM}$ $57.29\pm 5.32$
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $70.42\pm 2.36$
Table 4: Pretraining ++ fine-tuning Llama-3.2-1B-Instruct accuracy on pretraining dataset paraphrase and fine-tuning dataset rotten_tomatoes and yelp by Next Token Prediction ($\mathcal{L}_{\rm LLM}$) loss vs. $\mathcal{L}_{\rm LLM-JEPA}$ loss (our method). Note that $\mathcal{L}_{\rm LLM-JEPA}$ is applied only at pretraining. We tune $lr_{pre}$ and $lr_{ft}$ by $\mathcal{L}_{\rm LLM}$, and stick to them in LLM-JEPA pretraining. We run pretraining 5 times, and for each pretrained model, we run fine-tuning 5 times. Average accuracy and standard deviation are reported. We also report $p$-value of paired, single-tailed $t$-Test.
FT Dataset Method Accuracy (%) $\uparrow$ $p$-value $\downarrow$ Config
rotten_tomatoes $\mathcal{L}_{\rm LLM}$ $56.57\pm 1.66$ 7.38e47.38e-4 $lr_{pre}=8e-5,lr_{ft}=4e-5$
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $57.76\pm 1.33$ $\lambda=0.5,k=2$, same $lr_{pre},lr_{ft}$
yelp $\mathcal{L}_{\rm LLM}$ $26.46\pm 0.92$ 1.00e31.00e-3 $lr_{pre}=8e-5,lr_{ft}=8e-5$
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $27.15\pm 0.93$ $\lambda=0.5,k=2$, same $lr_{pre},lr_{ft}$
A.2 Ablation Study on the Role of $\mathcal{L}_{\rm LLM}$
One limitation of eq.˜2 is that the contribution of $\mathcal{L}_{\rm LLM}$ cannot be effectively reduced to 0. To address this, we introduce an additional hyperparameter $\gamma$ to explicitly control its relative strength:
$\mathcal{L}_{\rm LLM-JEPA}=\underbrace{\gamma\times\sum_{\ell=2}^{L}\mathcal{L}_{\rm LLM}({\rm Text}_{1:\ell-1},{\rm Text}_{\ell})}_{\text{generative capabilities}}+\lambda\times\underbrace{d({\rm Pred}({\rm Enc}({\rm Text})),{\rm Enc}({\rm Code}))}_{\text{abstraction capabilities}},$ (3)
We vary the ratio $\gamma/\lambda$ within $[0,1]$ while enforcing $\max(\gamma,\lambda)=1$ to maintain a constant learning rate. Table˜5 shows that $\mathcal{L}_{\rm LLM}$ remains essential for generative performance: when $\gamma=0$, the fine-tuned model produces only empty outputs. This indicates that the JEPA component primarily serves as a regularization term, complementing the generative loss.
Table 5: Fine-tuning accuracy on dataset NL-RX-SYNTH with $\mathcal{L}_{\rm LLM-JEPA}$ loss (ours) over various $\gamma/\lambda$. Configuration is $lr=2e-5,\lambda=1,k=0$. We maintain $\max(\gamma,\lambda)=1.0$ to use a fixed lrlr. Each cell runs five times. Average accuracy and standard deviation are reported. When $\gamma=0.0$, it generate only empty output.
$\gamma/\lambda$ Config Accuracy (%) $\uparrow$
$0.0$ $\gamma=0.0,\lambda=1.0$ $0.00\pm 0.00$
$0.001$ $\gamma=0.01,\lambda=1.0$ $1.38\pm 0.06$
$0.1$ $\gamma=0.1,\lambda=1.0$ $45.80\pm 5.04$
$1.0$ $\gamma=1.0,\lambda=1.0$ $70.42\pm 2.36$
$10.0$ $\gamma=1.0,\lambda=0.1$ $67.52\pm 1.45$
$100.0$ $\gamma=1.0,\lambda=0.01$ $66.83\pm 3.89$
$\infty$ $\gamma=1.0,\lambda=0.0$ $57.29\pm 5.32$
A.3 Hyperparameter Tuning for LLM-JEPA
Despite its strong accuracy gains, LLM-JEPA introduces two additional hyperparameters. As shown in fig.˜3, the optimal configuration may occur at any point in the grid $(\lambda,k)\in\{0.5,1.0,2.0,4.0\}\times\{0,1,2,3,4\}$, which imposes a significant cost for hyperparameter tuning. While we have not identified an efficient method to explore this space, we empirically observe that adjacent grid points often yield similar accuracy, suggesting the potential for a more efficient tuning algorithm.
(a) Llama on GSM8K, lr=2e5lr=2e-5
(b) Llama on Spider, lr=1e5lr=1e-5
(c) Gemma on SYNTH, lr=1e5lr=1e-5
(d) OpenELM on SYNTH, lr=8e5lr=8e-5
(e) OLMo on SYNTH, lr=8e5lr=8e-5
(f) Llama on SYNTH, Pretrain, lr=8e5lr=8e-5
Figure 3: In general we didn’t find any pattern on where the best accuracy could appear. It could be at either high-end or low-end of either $\lambda$ or $k$. Furthermore, there can be dips and spikes in random locations. Nonetheless, adjacent cells have close accuracy most of times, and sweeping $(k,\lambda)\in\{0,1,2,3,4\}\times\{0.5,1,2,4\}$ normally yield satisfiable results. Each cell is an average of five runs, $epoch=4$.
A.4 Additional Generation Examples
Table˜7 presents additional examples generated by fine-tuning Llama-3.2-1B-Instruct on the NL-RX-SYNTH dataset using $\mathcal{L}_{\rm LLM}$ and $\mathcal{L}_{\rm LLM-JEPA}$, respectively.
Table 6: Regular expressions generated by Llama-3.2-1B-Instruct after fine-tuning with $\mathcal{L}_{\rm LLM}$ loss and $\mathcal{L}_{\rm LLM-JEPA}$ loss (ours). Color code: wrong, extra, missing
Ground Truth $\mathcal{L}_{\rm LLM}$ $\mathcal{L}_{\rm LLM-JEPA}$ (ours)
lines not having the string ’dog’ followed by a number, 3 or more times
 ((dog.*[0-9].*)3,)  ((dog.*[0-9].*)3,)  ((dog.*[0-9].*){3,})
lines containing ending with a vowel, zero or more times
.*(.*)(([AEIOUaeiou])*).* (.*)(([AEIOUaeiou])*)* (.*)(.*){([AEIOUaeiou])*}
lines with a number or a character before a vowel
(([0-9])|(.)).*([AEIOUaeiou]).* (([0-9])|(.)).*([AEIOUaeiou]).*.* (([0-9])|(.)).*([AEIOUaeiou]).*
lines with words with the string ’dog’, a letter, and a number
((([0-9])&(dog))|([A-Za-z]))* ((([0-9])&(dog))|([A-Za-z]))* (([0-9])&(dog))|(([A-Za-z])*)
[…]
Table 7: More regular expressions generated by Llama-3.2-1B-Instruct after fine-tuning with $\mathcal{L}_{\rm LLM}$ loss and $\mathcal{L}_{\rm LLM-JEPA}$ loss (ours). Color code: wrong, extra, missing
Ground Truth $\mathcal{L}_{\rm LLM}$ $\mathcal{L}_{\rm LLM-JEPA}$ (ours)
lines ending with a vowel or starting with a character
([AEIOUaeiou].*[A-Za-z].*)+ ([AEIOUaeiou].*[A-Za-z].*)+ ([AEIOUaeiou].*[A-Za-z].*)+
ines containing either a lower-case letter, a vowel, or a letter
((.*)([AEIOUaeiou]))|((.)(.*)) (.*)(([AEIOUaeiou])|((.)(.*))) (.*)(([AEIOUaeiou])|((.)(.*)))
lines starting with the string ’dog’ before a vowel
(([A-Za-z])7,).*(dog).* (([A-Za-z])7,).*(dog).*.* (([A-Za-z])7,).*(dog).*
lines not containing a letter and the string ’dog’
((([A-Z])+)|([a-z]))(.*) ((([A-Z])+)|([a-z]))(.*)+ ((([A-Z])+)|([a-z]))(.*)
lines with a character before a vowel and the string ’dog’, zero or more times
.*(.)&([0-9])&(dog).* .*(.)&([0-9])&(dog).*.* .*(.)&([0-9])&(dog).*.*.*
lines with a vowel at least once before not a character
(([A-Za-z])+).*(~([0-9])).* (([A-Za-z])+).*(~([0-9])).*.* (([A-Za-z])+).*(~([0-9])).*
A.5 Overfitting Behavior in LoRA Fine-Tuning
We also conducted experiments to examine whether LoRA fine-tuning with $\mathcal{L}_{\rm LLM\text{-}JEPA}$ exhibits similar resistance to overfitting. As shown in fig.˜5, accuracy under $\mathcal{L}_{\rm LLM\text{-}JEPA}$ generally continues to improve with additional epochs, whereas fine-tuning with $\mathcal{L}_{\rm LLM}$ shows clear signs of overfitting. Notably, the standard deviation is much higher than in full fine-tuning, likely reflecting the lower capacity of LoRA fine-tuning. An interesting pattern emerges: for $\mathcal{L}_{\rm LLM\text{-}JEPA}$, larger standard deviations often coincide with dips in accuracy, whereas for $\mathcal{L}_{\rm LLM}$ they tend to accompany accuracy spikes. This suggests that such fluctuations may be unreliable indicators of generalization quality.
Figure 5: LLM-JEPA resists overfitting in LoRA fine-tuning. Fine-tuning with $\mathcal{L}_{\rm LLM-JEPA}$ loss (our method) resists overfitting. When fine-tuning with $\mathcal{L}_{\rm LLM}$ loss start to overfit, $\mathcal{L}_{\rm LLM-JEPA}$ kept improving. However the trend is not as stable as in full fine-tuning, possibly due to limited capacity of LoRA fine-tuning.
Table 8: Fine-tuning accuracy on dataset NL-RX-SYNTH by Next Token Prediction ($\mathcal{L}_{\rm LLM}$) loss vs. $\mathcal{L}_{\rm LLM-JEPA}$ loss (our method). Each cell is the best possible accuracy over a set of configurations. Each configuration runs five times. Average accuracy and standard deviation are reported. We also report $p$-value of paired, single-tailed $t$-Test.
Model Method Accuracy (%) $\uparrow$ $p$-value $\downarrow$ Config
gemma-2-2b-it $\mathcal{L}_{\rm LLM}$ $33.65\pm 3.24$ 5.5e35.5e-3 lr=1e5lr=1e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $43.12\pm 2.61$ $\lambda=2,k=4$, same lrlr
OpenELM-1_1B-Instruct $\mathcal{L}_{\rm LLM}$ $12.07\pm 1.81$ 5.1e45.1e-4 lr=8e5lr=8e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $25.40\pm 2.40$ $\lambda=4,k=3$, same lrlr
OLMo-2-0425-1B-Instruct $\mathcal{L}_{\rm LLM}$ $87.09\pm 0.36$ 2.5e32.5e-3 lr=8e5lr=8e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $87.52\pm 0.29$ $\lambda=2,k=0$, same lrlr
Table 9: Fine-tuning accuracy by model Llama-3.2-1B-Instruct, $\mathcal{L}_{\rm LLM}$ loss vs. $\mathcal{L}_{\rm LLM-JEPA}$ loss (our method). Each cell is the best possible accuracy over a set of configurations. Each configuration runs five times. Average accuracy and standard deviation are reported. We also report $p$-value of paired, single-tailed $t$-Test.
Dataset Method Accuracy (%) $\uparrow$ $p$-value $\downarrow$ Config
NL-RX-SYNTH $\mathcal{L}_{\rm LLM}$ $57.29\pm 5.32$ 1.0e31.0e-3 lr=2e5lr=2e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $71.46\pm 1.34$ $\lambda=1,k=1$, same lrlr
NL-RX-TURK $\mathcal{L}_{\rm LLM}$ $22.49\pm 1.91$ 2.4e42.4e-4 lr=2e5lr=2e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $30.94\pm 1.13$ $\lambda=1,k=1$, same lrlr
GSM8K $\mathcal{L}_{\rm LLM}$ $32.36\pm 0.58$ 9.6e59.6e-5 lr=2e5lr=2e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $36.36\pm 0.20$ $\lambda=0.5,k=4$, same lrlr
Spider $\mathcal{L}_{\rm LLM}$ $47.52\pm 2.44$ 4.0e34.0e-3 lr=4e5lr=4e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $50.55\pm 2.08$ $\lambda=1,k=3$, same lrlr
A.6 Structured Representations Induced by LLM-JEPA
We also examine the representation space to better understand how LLM-JEPA regularizes learned features. Specifically, we plot $t$-SNE embeddings for both $\operatorname{Text}$ and $\operatorname{Code}$ across three settings: the base model, a model fine-tuned with $\mathcal{L}_{\rm LLM}$, and a model fine-tuned with $\mathcal{L}_{\rm LLM\text{-}JEPA}$. As shown in fig.˜6, clear structure emerges after fine-tuning with $\mathcal{L}_{\rm LLM\text{-}JEPA}$. We hypothesize that $\mathcal{L}_{\rm LLM\text{-}JEPA}$ enforces structure in the representation space by constraining the mapping from $\operatorname{Enc}(\operatorname{Text})$ to $\operatorname{Enc}(\operatorname{Code})$ within a narrow subspace. If this is the case, the SVD decomposition of $\operatorname{Enc}(\operatorname{Text})-\operatorname{Enc}(\operatorname{Code})$ should yield significantly smaller singular values, which is confirmed in fig.˜7. Furthermore, we hypothesize that the mapping is approximately linear. To test this, we compute the least-squares regression error, and table˜10 supports this hypothesis. Together, these results suggest that LLM-JEPA promotes a near-linear transformation between $\operatorname{Text}$ and $\operatorname{Code}$ representations, which may underlie its accuracy improvements.
Table 10: LLM-JEPA is almost a linear transformation from $\operatorname{Enc}(\operatorname{Text})$ to $\operatorname{Enc}(\operatorname{Code})$.
[…]
Base model $3953.11$ $310.73$
$\mathcal{L}_{\rm LLM}$ $3035.01$ $341.80$
LLM-JEPA (Ours) $k=1$ $4.47$ $94.84$
LLM-JEPA (Ours) $k=0$ $4.04$ $16.82$
(a) Base model: No fine-tuning
(b) Baseline: Fine-tuned by NTP loss
(c) LLM-JEPA (Ours) $k=0$
(d) LLM-JEPA (Ours) $k=1$
Figure 6: $t$-SNE plot of Text and Code representations in (a) Base mode without fine-tuning, (b) Baseline that is fine-tuned with NTP loss, (c) LLM-JEPA (ours) with $k=0$, and (d) LLM-JEPA (ours) with $k=1$. Clearly LLM-JEPA (ours) induced nice structure on the representations while fine-tuning with NTP loss disrupted the structure in the base model.
Figure 7: The top 100 singular values of $\operatorname{Enc}(\operatorname{Text})-\operatorname{Enc}(\operatorname{Code})$. The curves of LLM-JEPA (ours) are a few magnitudes lower than that of base model and regular fine-tuning, meaning the mapping from Text to Code are confined within a narrow subspace, fostering the nice structure we see in Figure 6
A.7 Performance Across Model Sizes
We also evaluate LLM-JEPA across different model sizes. As shown in table˜11, we observe statistically significant improvements at all scales. Since there is no official 8B version of Llama-3.2, we instead use Llama-3.1-8B-Instruct, where performance collapsed due to the model’s difficulty in properly terminating regular expressions. To address this, we additionally evaluate using a startswith criterion—that is, a prediction is considered correct if the generated regular expression begins with the ground-truth expression, removing the need for exact termination. Under this metric, we again observe statistically significant accuracy improvements.
Table 11: Fine-tuning accuracy on NL-RX-SYNTH by Next Token Prediction ($\mathcal{L}_{\rm LLM}$) loss vs. $\mathcal{L}_{\rm LLM-JEPA}$ loss (our method). Each case runs five times. Average accuracy and standard deviation are reported. We also report $p$-value of paired, single-tailed $t$-Test. Note that Llama does not have official 3.2-8B, and we have to use 3.1-8B, which has a lower accuracy. Still LLM-JEPA sees significant improvement. We also evaluated on OLMo-2-7B.
Model Method Accuracy (%) $\uparrow$ $p$-value $\downarrow$ Config
Llama-3.2-1B-Instruct $\mathcal{L}_{\rm LLM}$ $57.29\pm 5.32$ 1.0e31.0e-3 lr=2e5lr=2e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $71.46\pm 1.34$ $\lambda=1,k=1$, same lrlr
Llama-3.2-3B-Instruct $\mathcal{L}_{\rm LLM}$ $74.55\pm 3.58$ $0.0352$ lr=2e5lr=2e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $77.16\pm 3.66$ $\lambda=2,k=0$, same lrlr
Llama-3.1-8B-Instruct $\mathcal{L}_{\rm LLM}$ $35.77\pm 6.60$ $0.0131$ lr=2e5lr=2e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $63.57\pm 16.81$ $\lambda=2.0,k=0$, same lrlr
OLMo-2-1124-7B-Instruct $\mathcal{L}_{\rm LLM}$ $87.26\pm 0.27$ $0.0345$ lr=2e5lr=2e-5
$\mathcal{L}_{\rm LLM-JEPA}$ (ours) $87.75\pm 0.33$ $\lambda=20,k=2$, same lrlr